OpenStudy (anonymous):

Verify the identity tan (x +pi/2) = -cot x

OpenStudy (anonymous):

I've gotten this far tan (pi/2 +×)= sin (pi/2 +×)/ cos ( pi/2 +x)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you need to use the "addition angle" formula for tangent

OpenStudy (anonymous):

@misty1212 hey! that's quite a nostalgic formula

OpenStudy (misty1212):

or use the addition angle formula for sine and cosine if you want to do what you wrote in the first step either way

OpenStudy (anonymous):

or you can, \[\sin(\frac{\pi}{2}+x)=\cos(x)\] \[\cos(\frac{\pi}{2}+x)=-\sin(x)\]

OpenStudy (anonymous):

sine, cosine and tangent formulae \[\sin(x+y)=sinxcosy+cosxsiny\]\[\cos(x+y)=cosxcosy-sinxsiny\] \[\tan(x+y)=\frac{tanx+tany}{1-tanxtany}\]

OpenStudy (anonymous):

OK that would give me tan (pi/2 +×) = sin (×) cos (pi/2) + cos (×) sin (pi/2) Right?

OpenStudy (anonymous):

I'm confused @ Nishshant_Garg and @misty1212

OpenStudy (anonymous):

@Nishant_Garg

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

that's not right allydiaz

OpenStudy (anonymous):

\[\tan(\frac{\pi}{2}+x)=\frac{\sin(\frac{\pi}{2}+x)}{\cos(\frac{\pi}{2}+x)}\] \[\frac{\sin(\frac{\pi}{2})cosx+\cos(\frac{\pi}{2}sinx)}{\cos(\frac{\pi}{2})cosx-\sin(\frac{\pi}{2})sinx}\]

OpenStudy (anonymous):

sorry that sinx is not inside cos on the numerator

OpenStudy (anonymous):

\[\sin(\frac{\pi}{2})=1\]\[\cos(\frac{\pi}{2})=0\]

OpenStudy (anonymous):

OK so how would it cancel out to leave behind -cos/sin?

OpenStudy (anonymous):

\[\frac{1 \times cosx+0 \times sinx}{0 \times cosx-1 \times sinx}=\frac{cosx}{-sinx}=-cotx\]

OpenStudy (anonymous):

you replace sin pi/2 with 1 and cos pi/2 with 0

OpenStudy (anonymous):

So cosx × sinx = cos x?